# Multi-Fidelity Bayesian Optimization With Across-Task Transferable Max-Value Entropy Search

### v2

@@ Abstract
In many applications, ranging from logistics to engineering, a designer is faced with a sequence of optimization tasks for which the objectives are in the form of black-box functions that are costly to evaluate. For example, the designer may need to tune the hyperparameters of neural network models for different learning tasks over time. Rather than evaluating the objective function for each candidate solution, the designer may have access to approximations of the objective functions, for which higher-fidelity evaluations entail a larger cost. Existing multi-fidelity black-box optimization strategies select candidate solutions and fidelity levels with the goal of maximizing the information accrued about the optimal value or solution for the current task. Assuming that successive optimization tasks are related, this paper introduces a novel information-theoretic acquisition function that balances the need to acquire information about the current task with the goal of collecting information transferable to future tasks. The proposed method includes shared inter-task latent variables, which are transferred across tasks by implementing particle-based variational Bayesian updates. Experimental results across synthetic and real-world examples reveal that the proposed provident acquisition strategy that caters to future tasks can significantly improve the optimization efficiency as soon as a sufficient number of tasks is processed.
@@ I-A Context and Scope
Numerous problems in logistics, science, and engineering can be formulated as black-box optimization tasks, in which the objective is costly to evaluate. Examples include hyperparameter optimization for machine learning [1], malware detection [2], antenna design [3], text to speech adaptation [4], material discovery [5], and resource allocation in wireless communication systems [6, 7, 8, 9]. To mitigate the problem of evaluating a costly objective function for each candidate solutions, the designer may have access to cheaper approximations of the optimization target. For example, the designer may be able to simulate a physical system using a digital twin that offers a controllable trade-off between cost and fidelity of the approximation [10, 11, 12, 13]. As shown in Fig. 1, higher-fidelity evaluations of the objective functions generally entail a larger cost, and the main challenge for the designer is to select a sequence of candidate solutions and fidelity levels that obtains the best solution within the available cost budget.
@@ I-B Related Work
Prior works developed MFBO by building on standard BO acquisition functions, including expected improvement (EI) in [24], upper confidence bound (UCB) in [25], and knowledge gradient (KG) in [23, 26]. EI-based MFBO does not account for the level of uncertainty in the surrogate model. This issue is addressed by UCB-based approaches, which, however, require a carefully selected parameter to balance exploitation and exploration. Finally, although KG-based methods achieve efficient global optimization without hyperparameters in the acquisition function, empirical studies in [21] show that they incur a high computational overhead.
@@ I-C Main Contributions
Assuming that successive optimization tasks are related, this paper introduces a novel information-theoretic acquisition function that balances the need to acquire information about the current task with the goal of collecting information transferable to future tasks. The proposed method, referred to as multi-fidelity transferable max-value entropy search (MFT-MES), includes shared inter-task latent variables, which are transferred across tasks by implementing particle-based variational Bayesian updates.
@@
  • We introduce the MFT-MES, a novel black-box optimization scheme tailored for settings in which the designer is faced with a sequence of related optimization tasks. MFT-MES builds on MF-MES [21] by selecting candidate solutions and fidelity levels that maximize the information gain per unit cost. The information gain in MFT-MES accounts not only for the information about the optimal value of the current objective, as in MF-MES, but also for the information accrued on the inter-task shared latent parameters that can be transferred to future tasks. To this end, MFT-MES models the latent parameters as random quantities whose distributions are updated and transferred across tasks.
@@ I-D Organization
The rest of the paper is organized as follows. Sec. II formulates the sequential multi-task black-box optimization problem, and reviews the MF-GP surrogate model considered in the paper. Sec. III presents the baseline implementation of MF-MES, and illustrates the optimization over surrogate model parameters. The proposed MFT-MES method and the Bayesian update of the shared parameters are introduced in Sec. IV. Experimental results on synthetic optimization tasks and real-world applications are provided in Sec. V. Finally, Sec. VI concludes the paper.
@@ II-A Sequential Multi-Task Black-Box Optimization
We consider a setting in which optimization tasks, defined on a common input space $\mathcal{X}\subseteq\mathbbm{R}^{d}$, are addressed sequentially. Each $n$-th task, with $n=1,2,...$, consists of the optimization of a black-box expensive-to-evaluate objective function $f_{n}(\mathbf{x})$. Examples include the optimization of hyperparameters for machine learning models and experimental design [37, 38]. The objective functions $f_{n}(\mathbf{x})$ are assumed to be drawn according to a common parametric distribution $p(f(\mathbf{x})|\bm{\theta})$ in an independent identical distributed (i.i.d.) manner, i.e.,
@@
Furthermore, the parameter vector $\bm{\theta}$ identifying the distribution $p(f(\mathbf{x})|\bm{\theta})$ is unknown, and it is assigned a prior distribution $p(\bm{\theta})$, i.e.,
@@
Note that, by (1) and (2), the objective function $f_{1}(\mathbf{x}),f_{2}(\mathbf{x}),...$ are not independent, since having information on any function $f_{n}(\mathbf{x})$ would reduce uncertainty on the parameters $\bm{\theta}$, thus also providing information about other function $f_{n^{\prime}}(\mathbf{x})$ with $n^{\prime}\neq n$.
@@
In order to account for the cost of accessing the objective function $f_{n}(\mathbf{x})$, we follow the multi-fidelity formulation, whereby evaluating function $f_{n}(\mathbf{x})$ at some querying point $\mathbf{x}$ with fidelity level $m$ entails a cost $\lambda^{(m)}>0$ [39, 40]. Different fidelity levels may correspond to training processes with varying number of iterations for hyperparameters optimization, or to simulations of a physical process with varying levels of accuracy for experimental design.
@@ II-C Multi-fidelity Gaussian Process
Multi-fidelity GP (MFGP) provides a surrogate model for the objective functions $(f_{n}^{(1)}(\mathbf{x}),...,f_{n}^{(M)}(\mathbf{x}))$ across all $M$ fidelity levels [42, 43, 44]. This is done by defining a kernel function of the form $k_{\bm{\theta}}\big{(}(\mathbf{x},m),(\mathbf{x}^{\prime},m^{\prime})\big{)}$ that captures the correlations between the function values $f_{n}^{(m)}(\mathbf{x})$ and $f_{n}^{(m^{\prime})}(\mathbf{x}^{\prime})$ for any two inputs $\mathbf{x}$ and $\mathbf{x}^{\prime}$, and for any two fidelity levels $m$ and $m^{\prime}$. Examples of such kernels include the co-kriging model in [44] and the intrinsic coregionalization model (ICM) kernel [42, 45], which is expressed as
@@ III-A Multi-Fidelity Max-Value Entropy Search
where the set =$\mathcal{F}=\{f_{n,s}^{*}\}_{s=1}^{S}$ collects $S$ samples drawn from distribution $p_{\bm{\theta}}(f_{n}^{*}|\mathbf{x},\mathcal{D}_{n,t})$, which can be obtained via Gumbel sampling [19] and function $H_{\bm{\theta}}^{\text{MF-MES}}(\mathbf{x},m,f_{n}^{*})$ is defined in (22).
@@ IV-A Multi-Fidelity Transferable Max-Value Entropy Search
where the distribution $p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})$ is the posterior distribution on the shared model parameters $\bm{\theta}$. This distribution represents the transferable knowledge extracted from all the available observations at round $t$ for task $n$.
@@
where $\text{Var}(y_{n}^{(m)})$ represents the variance of random variable $y^{(m)}_{n}$ following the mixture of Gaussian distribution $\mathbb{E}_{p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})}[\mathcal{N}(y_{n}^{(m)}|\mu^{(m)}_{\bm{\theta}}(\mathbf{x}),\sigma_{\bm{\theta}}^{(m)2}(\mathbf{x}))]$. The tightness of upper bound (32) depends on the diversity of the Gaussian components with respect to the distribution $p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})$, becoming more accurate when the components of the mixture tend to the same Gaussian distribution.
@@ 21       Evaluate the function $\Omega(\bm{\theta}_{v}^{(r)})$ as in (37)
$(\mathbf{x}_{n,t+1},m_{n,t+1})=$ $\arg\max\limits_{\begin{subarray}{c}\mathbf{x}\in\mathcal{X}\\ m\in\mathcal{M}\end{subarray}}\frac{1}{\lambda^{(m)}}\Big{[}\alpha^{\text{MF-MES}}(\mathbf{x},m)$
$+\beta\Delta\alpha^{\text{MFT-MES}}(\mathbf{x},m)\Big{]},$ (35)
@@ IV-B Bayesian Learning for the Parameter Vector $\bm{\theta}$
In SVGD [34], the posterior distribution $p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})$ is approximated via a set of $V$ particles $\{\bm{\theta}_{1},\bm{\theta}_{2},...,\bm{\theta}_{V}\}$ that are iteratively transported to minimize the variational inference (VI) objective denoted by the Kullback–Leibler (KL) divergence $\mathrm{KL}(q(\bm{\theta})||p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t}))$ over a distribution $q(\bm{\theta})$ represented via particles $\{\bm{\theta}_{1},\bm{\theta}_{2},...,\bm{\theta}_{V}\}$. This is done via functional gradient descent in the reproducing kernel Hilbert space (RKHS).
@@
V Experiments

### v3

@@ Abstract
In many applications, ranging from logistics to engineering, a designer is faced with a sequence of optimization tasks for which the objectives are in the form of black-box functions that are costly to evaluate. Furthermore, higher-fidelity evaluations of the optimization objectives often entail a larger cost. Existing multi-fidelity black-box optimization strategies select candidate solutions and fidelity levels with the goal of maximizing the information about the optimal value or the optimal solution for the current task. Assuming that successive optimization tasks are related, this paper introduces a novel information-theoretic acquisition function that balances the need to acquire information about the current task with the goal of collecting information transferable to future tasks. The proposed method transfers across tasks distributions over parameters of a Gaussian process surrogate model by implementing particle-based variational Bayesian updates. Theoretical insights based on the analysis of the expected regret substantiate the benefits of acquiring transferable knowledge across tasks. Furthermore, experimental results across synthetic and real-world examples reveal that the proposed acquisition strategy that caters to future tasks can significantly improve the optimization efficiency as soon as a sufficient number of tasks is processed.
@@ I-A Context and Scope
Numerous problems in logistics, science, and engineering can be formulated as black-box optimization tasks, in which the objective is costly to evaluate. Examples include hyperparameter optimization for machine learning [1], malware detection [2], antenna design [3], text to speech adaptation [4], and resource allocation in wireless communication systems [5, 6, 7]. To mitigate the problem of evaluating a costly objective function for each candidate solutions, the designer may have access to cheaper approximations of the optimization target. For example, the designer may be able to simulate a physical system using a digital twin that offers a controllable trade-off between cost and fidelity of the approximation [8, 9, 10]. As shown in Fig. 1, higher-fidelity evaluations of the objective functions generally entail a larger cost, and the main challenge for the designer is to select a sequence of candidate solutions and fidelity levels that obtains the best solution within the available cost budget.
@@ I-B Related Work
Prior works developed MFBO by building on standard BO acquisition functions, including expected improvement (EI) in [21], upper confidence bound (UCB) in [22], and knowledge gradient (KG) in [20]. EI-based MFBO does not account for the level of uncertainty in the surrogate model. This issue is addressed by UCB-based approaches, which, however, require a carefully selected parameter to balance exploitation and exploration. Finally, although KG-based methods achieve efficient global optimization without hyperparameters in the acquisition function, empirical studies in [18] show that they incur a high computational overhead.
@@ I-C Main Contributions
Assuming that successive optimization tasks are related, this paper introduces a novel information-theoretic acquisition function that balances the need to acquire information about the current task with the goal of collecting information transferable to future tasks. The proposed method, referred to as multi-fidelity transferable max-value entropy search (MFT-MES), includes shared GP model parameters, which are transferred across tasks by implementing particle-based variational Bayesian updates.
@@
  • We introduce the MFT-MES, a novel black-box optimization scheme tailored for settings in which the designer is faced with a sequence of related optimization tasks. MFT-MES builds on MF-MES [18] by selecting candidate solutions and fidelity levels that maximize the information gain per unit cost. The information gain in MFT-MES accounts not only for the information about the optimal value of the current objective, as in MF-MES, but also for the information accrued on the shared GP model parameters that can be transferred to future tasks. To this end, MFT-MES models the latent parameters as random quantities whose distributions are updated and transferred across tasks.
@@
  • We provide an analysis of the expected regret bound for a stylized version of MFT-MES. The theoretical results substantiate the benefits of exploring the candidate solutions that are likely to provide knowledge transferable across tasks.
@@
This work was partially presented as a conference paper in [33]. The conference version presents a brief presentation of the proposed acquisition strategy, along with limited experiments. In contrast, this paper provides full details on the proposed MFT-MES scheme, along with a theoretical analysis of the expected regret bound, as well as with comprehensive experimental results.
@@ I-D Organization
The rest of the paper is organized as follows. Sec. II formulates the sequential multi-task black-box optimization problem, and reviews the MF-GP surrogate model considered in the paper. Sec. III presents the baseline implementation of MF-MES, and illustrates the optimization over surrogate model parameters. The proposed MFT-MES method and the Bayesian update of the shared parameters are introduced in Sec. IV. A theoretical analysis of MFT-MES is provided in Sec. V. Experimental results on synthetic optimization tasks and real-world applications are provided in Sec. VI. Finally, Sec. VII concludes the paper.
@@ II-A Sequential Multi-Task Black-Box Optimization
We consider a setting in which optimization tasks, defined on a common input space $\mathcal{X}\subseteq\mathbbm{R}^{d}$, are addressed sequentially. Each $n$-th task, with $n=1,2,...$, consists of the optimization of a black-box expensive-to-evaluate objective function $f_{n}(\mathbf{x})$. Examples include the optimization of hyperparameters for machine learning models and experimental design [34]. The objective functions $f_{n}(\mathbf{x})$ are assumed to be drawn according to a parametric stochastic process $\mathcal{P}_{\bm{\theta}}(f(\mathbf{x}))$ in an independent identical distributed (i.i.d.) manner, i.e.,
@@
Furthermore, the parameter vector $\bm{\theta}$ identifying the stochastic process $\mathcal{P}_{\bm{\theta}}(f(\mathbf{x}))$ is unknown, and it is assigned a prior distribution $p(\bm{\theta})$, i.e.,
@@
Note that, by assumptions (1) and (2), the objective functions $f_{1}(\mathbf{x}),f_{2}(\mathbf{x}),...$ are not independent unless one conditions on parameter vector $\bm{\theta}$. In fact, having information on any function $f_{n}(\mathbf{x})$ reduces uncertainty on the parameter vector $\bm{\theta}$, thus providing information about other functions $f_{n^{\prime}}(\mathbf{x})$ with $n^{\prime}\neq n$.
@@
In order to account for the cost of accessing the objective function $f_{n}(\mathbf{x})$, we follow the multi-fidelity formulation, whereby evaluating function $f_{n}(\mathbf{x})$ at some querying point $\mathbf{x}$ with fidelity level $m$ entails a cost $\lambda^{(m)}>0$ [35]. Different fidelity levels may correspond to training processes with varying number of iterations for hyperparameters optimization, or to simulations of a physical process with varying levels of accuracy for experimental design.
@@ II-C Multi-fidelity Gaussian Process
Multi-fidelity GP (MFGP) provides a surrogate model for the objective functions $(f_{n}^{(1)}(\mathbf{x}),...,f_{n}^{(M)}(\mathbf{x}))$ across all $M$ fidelity levels [37, 38]. This is done by defining a kernel function of the form $k_{\bm{\theta}}\big{(}(\mathbf{x},m),(\mathbf{x}^{\prime},m^{\prime})\big{)}$ that captures the correlations between the function values $f_{n}^{(m)}(\mathbf{x})$ and $f_{n}^{(m^{\prime})}(\mathbf{x}^{\prime})$ for any two inputs $\mathbf{x}$ and $\mathbf{x}^{\prime}$, and for any two fidelity levels $m$ and $m^{\prime}$. Examples of such kernels include the co-kriging model in [38] and the intrinsic coregionalization model (ICM) kernel [37, 39], which is expressed as
@@ III-A Multi-Fidelity Max-Value Entropy Search
where the set =$\mathcal{F}=\{f_{n,s}^{*}\}_{s=1}^{S}$ collects $S$ samples drawn from distribution $p_{\bm{\theta}}(f_{n}^{*}|\mathbf{x},\mathcal{D}_{n,t})$, which can be obtained via Gumbel sampling (see [16, Section 3.1] for details) and function $H_{\bm{\theta}}^{\text{MF-MES}}(\mathbf{x},m,f_{n}^{*})$ is defined in (22).
@@ IV-A Multi-Fidelity Transferable Max-Value Entropy Search
where the distribution $p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})$ is the posterior distribution on the shared model parameters $\bm{\theta}$. This posterior distribution represents the knowledge extracted from all the observations available at round $t$ of task $n$ that can be transferred to future tasks. It is obtained as
$p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})=\frac{p(\bm{\theta},\mathcal{D}_{n-1},\mathcal{D}_{n,t})}{p(\mathcal{D}_{n-1},\mathcal{D}_{n,t})}$ (30)
with the joint distribution
$p(\bm{\theta},\mathcal{D}_{n-1},\mathcal{D}_{n,t})$
$=p(\bm{\theta})\cdot\prod_{i=1}^{n-1}p(\mathcal{D}_{i}|\bm{\theta})p(\mathcal{D}_{n,t}|\bm{\theta})$
$=p(\bm{\theta})\cdot\prod_{i=1}^{n-1}p(\mathbf{y}_{i}|\mathbf{x},\bm{\theta})p(\mathbf{x})\cdot p(\mathbf{y}_{n,t}|\mathbf{x},\bm{\theta})p(\mathbf{x}),$ (31)
[…]
@@
where $\text{Var}(y_{n}^{(m)})$ represents the variance of random variable $y^{(m)}_{n}$ following the mixture of Gaussian distribution $\mathbb{E}_{p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})}[\mathcal{N}(y_{n}^{(m)}|\mu^{(m)}_{\bm{\theta}}(\mathbf{x}),\sigma_{\bm{\theta}}^{(m)2}(\mathbf{x}))]$. The tightness of upper bound (34) depends on the diversity of the Gaussian components with respect to the distribution $p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})$, becoming more accurate when the components of the mixture tend to the same Gaussian distribution [44].
@@ 21       Evaluate the function $\Omega(\bm{\theta}_{v}^{(r)})$ as in (39)
$(\mathbf{x}_{n,t+1},m_{n,t+1})=$ $\arg\max\limits_{\begin{subarray}{c}\mathbf{x}\in\mathcal{X}\\ m\in\mathcal{M}\end{subarray}}\mathbbm{E}_{p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})}[\alpha_{\bm{\theta}}^{\text{MF-MES}}(\mathbf{x},m)]$
$+\beta\cdot\frac{\Delta\alpha^{\text{MFT-MES}}(\mathbf{x},m)}{\lambda^{(m)}},$ (37)
@@ IV-B Bayesian Learning for the Parameter Vector $\bm{\theta}$
In SVGD [30], the posterior distribution $p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})$ is approximated via a set of $V$ particles $\{\bm{\theta}_{1},\bm{\theta}_{2},...,\bm{\theta}_{V}\}$ that are iteratively transported to minimize the variational inference (VI) objective denoted by the Kullback–Leibler (KL) divergence $\mathrm{KL}(q(\bm{\theta})||p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t}))$ over a distribution $q(\bm{\theta})$ represented via particles $\{\bm{\theta}_{1},\bm{\theta}_{2},...,\bm{\theta}_{V}\}$. This is done via functional gradient descent in the reproducing kernel Hilbert space (RKHS). Under mild assumptions on the kernel derivatives and log-loss gradient in (39), SVGD using $V$ particles converges to the true posterior distribution at rate $\mathcal{O}(1/\sqrt{V})$ [46, Theorem 1].
@@
V Theoretical Analysis
In this section, we establish an upper bound on the expected regret for a simplified version of MFT-MES by building on the expected regret bound presented in [16] for MES.
Throughout, we assume as in [16], that the objective functions $f_{n}(\mathbf{x})$ follow the GP (10) with a given ground-truth parameter vector $\bm{\theta}^{*}$. Moreover, for each task $n$, we evaluate the performance at time $t$ via average regret
$\bar{r}_{n,t}(\mathbf{x}_{n,t}|\bm{\theta}^{*})=\mathbbm{E}[f_{n}^{*}-f_{n}(\mathbf{x}_{n,t})|\mathcal{D}_{n,t-1}],$ (40)
where the expectation is taken with respect to the posterior distribution (18) with $\bm{\theta}=\bm{\theta}^{*}$. The cumulative expected regret up to time $T$ for a task $n$ is accordingly defined as
$\mathcal{R}_{n,T}(\mathbf{X}_{n,T}|\bm{\theta}^{*})=\sum_{t=1}^{T}\bar{r}_{n,t}(\mathbf{x}_{n,t}|\bm{\theta}^{*}).$ (41)
V-A Two-Phase MFT-MES
To facilitate the analysis of the proposed MFT-MES scheme, we study a generally less efficient variant that retains the essential features of the approach. Fixing the fidelity level $M$, the simplified MFT-MES protocol under study splits the iterations for each task $n$ into two separate phases, one devoted to per-task exploration via MES, and the other dedicated to the acquisition of information about the parameter vector $\bm{\theta}^{*}$. Note that MFT-MES in Algorithm 2 adopts a more flexible strategy that integrates these two phases.
To elaborate, for each task $n$, we divide the total number of iterations $T$ into two parts via the selection of a hyperparameter $c\in[0,1]$. For the first $\lceil c\cdot T\rceil$ iterations, the next candidate solution is selected using MES as
$\mathbf{x}_{n,t+1}=\arg\max\limits_{\mathbf{x}\in\mathcal{X}}\mathbb{E}_{\bm{\theta}\sim p(\bm{\theta}|\mathcal{D}_{n-1})}\big{[}\bar{\alpha}_{\bm{\theta}}^{\text{MES}}(\mathbf{x})\big{]}$ (42)
for $t+1\leq c\cdot T$, where $\bar{\alpha}_{\bm{\theta}}^{\text{MES}}(\mathbf{x})=g(\frac{\hat{m}_{n,t}(\bm{\theta}^{*})-\mu_{\bm{\theta}^{*}}(\mathbf{x})}{\sigma_{\bm{\theta}^{*}}(\mathbf{x})})$ is the acquisition function of MES with $\hat{m}_{n,t}(\bm{\theta}^{*})=\mathbb{E}_{f_{n}(\mathbf{x})}[f_{n}^{*}]$ being the expected maximum value of the function [16]. Note that, unlike the original MES scheme, the acquisition procedure (42) applies an average over the current posterior distribution $p(\bm{\theta}|\mathcal{D}_{n-1})$ for the parameter vector $\bm{\theta}$, given that the ground-truth $\bm{\theta}^{*}$ is unknown.
In the second set of $T-\lceil c\cdot T\rceil$ iterations, the optimizer attempts to maximize the information gain about parameter vector $\bm{\theta}^{*}$ by selecting the batch of solutions $\mathbf{X}=\{\mathbf{x}_{n,t}\}_{t=\lceil c\cdot T\rceil+1}^{T}$ which maximize the mutual information
$\{\mathbf{x}_{n,t}\}_{t=\lceil c\cdot T\rceil+1}^{T}=\arg\max\limits_{\mathbf{X}\subset\mathcal{X}}I(\bm{\theta};\mathbf{y}|\mathbf{X}),$ (43)
where $\mathbf{y}$ is the observation vector of size $T-\lceil c\cdot T\rceil$ corresponding to the candidates $\mathbf{X}$. Note that, the objective (43) simplifies (33) in MFT-MES by (i) selecting the $T-\lceil c\cdot T\rceil$ candidates as a batch, rather than sequentially; and (ii) disregarding the data $\mathcal{D}_{n-1}$ collected for previous tasks.
V-B Technical Assumptions
To describe the technical assumptions on which the analysis is based, we first put forth two definitions. These will be used to account for the gap between MES with known parameter vector $\bm{\theta}^{*}$ and the acquisition function (42), which relies on the posterior distribution $p(\bm{\theta}|\mathcal{D}_{n-1})$.
Definition 1 ($\varepsilon$-sharpness).
For a random parameter vector $\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},\Sigma)$, a function $h(\mathbf{x}|\bm{\theta})$ is said to be $\varepsilon$-sharp around $\bm{\theta}^{*}$ if the inequality
$\Big{|}\Big{|}\arg\min_{\mathbf{x}\in\mathcal{X}}\mathbb{E}_{\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},\Sigma)}\big{[}h(\mathbf{x}|\bm{\theta})\big{]}-\arg\min_{\mathbf{x}\in\mathcal{X}}h(\mathbf{x}|\bm{\theta}^{*})\Big{|}\Big{|}$
$\leq\varepsilon\cdot\mathrm{Tr}(\Sigma)$ (44)
holds, where $\mathrm{Tr}(\cdot)$ is the trace operator.
The $\varepsilon$-sharpness property (44) describes how sensitive the minimizer of the function $h(\mathbf{x}|\bm{\theta}^{*})$ is with respect to the perturbations applied to the parameter vector $\bm{\theta}^{*}$ (see, e.g., [49] for related definitions).
Definition 2 ($\varepsilon$-stability).
Consider the monotonically decreasing function
$g(u)=u\cdot\frac{\phi(u)}{2\Phi(u)}-\log(\Phi(u)),$ (45)
as well as a random parameter vector $\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},\Sigma)$. The function $h(\mathbf{x}|\bm{\theta})$ is $\varepsilon$-stable around $\bm{\theta}^{*}$ if it satisfies the inequality
$\Big{|}\Big{|}\arg\max\limits_{\mathbf{x}\in\mathcal{X}}\mathbb{E}_{\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},\Sigma)}\big{[}g(h(\mathbf{x}|\bm{\theta}))\big{]}$
$-\arg\min\limits_{\mathbf{x}\in\mathcal{X}}\mathbb{E}_{\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},\Sigma)}\big{[}h(\mathbf{x}|\bm{\theta})\big{]}\Big{|}\Big{|}\leq\varepsilon\cdot\mathrm{Tr}(\Sigma).$ (46)
The $\varepsilon$-stability property (46) describes how sensitive the optimizer of the function $\mathbb{E}_{\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},\Sigma)}\big{[}g(h(\mathbf{x}|\bm{\theta}))\big{]}$ is with respect to the position of the expectation operator, i.e., applied before or after the monotonic function $g(\cdot)$ (see e.g., [50, 51] for related definitions).
Assumption 1 (Normalized regret conditions).
The normalized regret
$\tilde{r}_{n,t+1}(\mathbf{x}|\bm{\theta}^{*})=\frac{\bar{r}_{n,t+1}(\mathbf{x}|\bm{\theta}^{*})}{\sigma_{\bm{\theta}*}(\mathbf{x})}$ (47)
is $\varepsilon_{1}$-sharp around $\bm{\theta}^{*}$ as per Definition 1 as well as $\varepsilon_{2}$-stable around $\bm{\theta}^{*}$ as per Definition 2. Also, the normalized regret $\tilde{r}_{n,t}(\mathbf{x}|\bm{\theta}^{*})$ is $L$-Lipschitz continuous.
Assumption 2 (Convergence of the parameter vector posterior).
The posterior $p(\bm{\theta}|\mathcal{D}_{n-1})$ converges in distribution as
[…]
for a covariate matrix $\Sigma_{\bm{\theta}^{*}}$ satisfying the inequality
$\text{Tr}(\Sigma_{\bm{\theta}^{*}})\leq\frac{\xi}{(1-c)^{\alpha}\cdot T^{\alpha}}$ (49)
for some constants $0<\xi<\infty$ and $\alpha>0$.
Assumption 2 can be justified via the Bernstein–von Mises theorem [52], which guarantees convergence of the posterior $p(\bm{\theta}|\mathcal{D}_{n-1})$ to a Gaussian distribution $\mathcal{N}(\bm{\theta}^{*},F(\bm{\theta}^{*})^{-1}/n)$, where $F(\bm{\theta})$ is the Fisher information matrix of the likelihood $p(y|\mathbf{x},\bm{\theta})$.
In particular, Assumption 2 is verified as long as the Fisher information matrix has the property
$\text{Tr}(F(\bm{\theta}^{*})^{-1})=\mathcal{O}\bigg{(}\frac{1}{(1-c)^{\alpha}\cdot T^{\alpha}}\bigg{)}$ (50)
for some $\alpha>0$. For $\alpha=1$, the condition (50) corresponds to the additivity of Fisher information, which holds for i.i.d. observations $\mathbf{y}=\{f(\mathbf{x}_{t})+\epsilon_{t}\}_{t=\lceil c\cdot T\rceil+1}^{T}$ in each task. Given that the observations $\mathbf{y}$ are generally correlated, the scaling with $\alpha=1$ in (50) may not be satisfied. Indeed, as discussed in [53], the Fisher information can be sub-additive, yielding an exponent $\alpha<1$, for positively correlated observations, or super-additive, corresponding to an exponent $\alpha>1$, for negatively correlated observations.
V-C Regret Analysis
The cumulative expected regret bound of the two-phase MFT-MES protocol described in Sec. V-A, is bounded in the following theorem.
Theorem 1 (Cumulative Expected Regret Bound of MFT-MES).
Under Assumptions 1 and 2, denoting $\varepsilon=\varepsilon_{1}+\varepsilon_{2}$, the cumulative expected regret of the two-phase MFT-MES scheme satisfies the inequality
$\mathcal{R}_{n,cT}(\mathbf{X}_{n,cT}|\bm{\theta}^{*})$
$\leq\left(\nu_{n,t_{n}^{*}}+\frac{\varepsilon\cdot L\cdot\xi}{(n-1)\cdot(1-c)^{\alpha}\cdot T^{\alpha}}\right)\sqrt{\zeta\cdot c\cdot T\cdot\rho_{n,cT}},$ (51)
where $\nu_{n,t}=\min_{\mathbf{x}}\tilde{r}_{n,t}(\mathbf{x}|\bm{\theta}^{*})$ is the minimum normalized expected regret (47) at time $t$; $t_{n}^{*}=\arg\max_{t\in\{1,...,c\cdot T\}}\nu_{n,t}$ is the time at which $\nu_{n,t_{n}}$ is maximized; $0<\xi<\infty$ is a constant defined in Assumption 2; $\zeta$ is defined as $\zeta=2/\log(1+\sigma^{-2})$; and the maximum information gain $\rho_{n,cT}$ is
$\rho_{n,cT}=\max_{A\subset\mathcal{X}:|A|=cT}I_{n}(\mathbf{y}_{A};\bm{f}_{A}),$ (52)
which is evaluated as in [17, Eq. (7)].
Theorem 1, which is proven in Appendix, shows that the cumulative expected regret decreases as the number of tasks $n$ grows large, reflecting the benefits of knowledge transferred across tasks. In particular, devoting a larger fraction of time $(1-c)$ to collect information about the shared parameters $\bm{\theta}^{*}$ ensures a faster rate of decrease of the cumulative regret with $n$. On the flip side, choosing an excessively large value of $(1-c)$ can cause a increase in the cumulative regret, as per the second term in (1), owing to the insufficient time devoted to the exploration of the current objective function. As verified via numerical results in the following section, this result suggests that MFT-MES can be advantageous as soon as $n$ is large enough, as in this case one can choose a value of $c$ near 1, while still guaranteeing a vanishing contribution of the term $(\varepsilon\cdot L\cdot\xi)/((n-1)\cdot(1-c)^{\alpha}\cdot T^{\alpha})$ in (1).
VI Experiments